# Feature learning as alignment: a structural property of gradient descent in non-linear neural networks

### v3

@@ Abstract
Understanding the mechanisms through which neural networks extract statistics from input-label pairs through feature learning is one of the most important unsolved problems in supervised learning. Prior works demonstrated that the gram matrices of the weights (the neural feature matrices, NFM) and the average gradient outer products (AGOP) become correlated during training, in a statement known as the neural feature ansatz (NFA). Through the NFA, the authors introduce mapping with the AGOP as a general mechanism for neural feature learning. However, these works do not provide a theoretical explanation for this correlation or its origins. In this work, we further clarify the nature of this correlation, and explain its emergence. We show that this correlation is equivalent to alignment between the left singular structure of the weight matrices and the newly defined pre-activation tangent features at each layer. We further establish that the alignment is driven by the interaction of weight changes induced by SGD with the pre-activation features, and analyze the resulting dynamics analytically at early times in terms of simple statistics of the inputs and labels. Finally, motivated by the observation that the NFA is driven by this centered correlation, we introduce a simple optimization rule that dramatically increases the NFA correlations at any given layer and improves the quality of features learned.
**footnotetext: Correspondence to: Daniel Beaglehole (dbeaglehole@ucsd.edu)$\dagger$\$$\dagger$\$footnotetext: Work partially done as an intern at Google DeepMind.
@@ 1 Introduction
Neural networks have emerged as the state-of-the-art machine learning methods for seemingly complex tasks, such as language generation [1], image classification [2], and visual rendering [3]. The precise reasons why neural networks generalize well have been the subject of intensive exploration, beginning with the observation that standard generalization bounds from statistical learning theory fall short of explaining their performance [4].
@@
A promising line of work emerged in the form of the neural tangent kernel, connecting neural networks to kernels in the wide limit [5, 6]. However, subsequent research showed that the success of neural networks relies critically on aspects of learning which are absent in kernel approximations [7, 8, 9, 10, 11]. Other work showed that low width suffices for gradient descent to achieve arbitrarily small test error [12], further refuting the idea that extremely wide networks are necessary.
@@
Subsequently, the success of neural networks has been largely attributed to feature learning - the ability of neural networks to learn statistics, measurements, and representations of data which are useful for downstream tasks. However, the specific mechanism through which features are learned is an important unsolved problem in deep learning theory. A number of works have studied the abilities of neural networks to learn features in structured settings [13, 14, 15, 16, 17, 18, 19]. Some of that work proves strict separation in terms of sample complexity between neural networks trained with stochastic gradient descent and kernels [20].
@@
The work above studies simple structure, such as learning from low-rank data or functions that are hierarchical compositions of simple elements. Recent work makes a big step towards generalizing these assumptions by proposing the neural feature ansatz (NFA) [21, 22], a general structure that emerges in the weights of trained neural networks. The NFA states that the gram matrix of the weights at a given layer (known as the neural feature matrix or NFM) is aligned with the average gradient outer product (AGOP) of the network with respect to the input to that layer. In particular, the NFM and AGOP are highly correlated in all layers of trained neural networks of general architectures, including practical models such as VGG [23], vision transformers [24], and GPT-family models [1].
@@
A major missing element is an explanation for how and why the AGOP and NFM become correlated through training with gradient descent. In this paper, we precisely describe the emergence of this correlation. We establish that the NFA is equivalent to alignment between the left singular structure of the weight matrices and the uncentered covariance of the pre-activation tangent kernel (PTK) (Section 2) features. We then introduce the centered neural feature correlation (C-NFC) which isolates this alignment process. We show empirically that the C-NFC is close to its maximum value of $1$ at early times, and fully captures the NFA at late times for a variety of architectures (fully-connected, convolutional, and attention layers) over a diverse collection of datasets (Section 3). Our experiments suggest that the C-NFC drives the development of the NFA. Through this centering, we identify analytically how the NFA emerges as a structural property of gradient descent. We provide further evidence by showing that the dynamics of the C-NFC can be understood analytically in terms of the statistics of the data and labels at early times (Section 3), and can be manipulated theoretically (Section 4).
@@ 2.1 Preliminaries
One can define two objects associated with neural networks that capture learned structure. For a given layer $\ell$, the neural feature matrix (NFM) $F_{\ell}$ is the gram matrix of the columns of the weight matrix $W^{(\ell)}$, i.e. $F_{\ell}\equiv(W^{(\ell)})^{\top}W^{(\ell)}$. $F_{\ell}$ depends on the right singular vectors (and corresponding singular values) of $W^{(\ell)}$. The second fundamental object we consider is the average gradient outer product (AGOP) $\bar{G}_{\ell}$, defined as $\bar{G}_{\ell}\equiv\frac{1}{n}\sum_{\alpha=1}^{n}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial x_{\ell}}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial x_{\ell}}^{\top}$.
@@
Ansatz 1 (Neural Feature Ansatz [21]).
@@
The relation between the NFM and the AGOP is significant, in part, because for a neural network that has learned enough information about the target function, the AGOP of this model with respect to the first-layer inputs will approximate the expected gradient outer product (EGOP) of the target function [25]. In particular, the EGOP of the target function contains task-specific structure that is completely independent of the model used to estimate it. Where the labels are generated from a particular target function $y(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$ on data sampled from a distribution $\mu$, the EGOP is defined as
@@
We verify for this task that the AGOP of the trained model resembles the EGOP (first and third panels of Figure 1). Here the NFA holds and the NFM (second panel) resembles the AGOP and therefore the EGOP as well. Therefore, the neural network has learned the model-independent and task-specific structure of the chain-monomial task in the right singular values and vectors of the first layer weight matrix, as these are determined by the NFM. In fact, previous works have demonstrated that the NFM of the first layer of a well-trained neural network is highly correlated with the AGOP of a fixed kernel method trained on the same dataset [21].
This insight has inspired iterative kernel methods which can match the performance of fully-connected networks [21, 26, 27] and improve over fixed convolutional kernels [22]. Additional prior works demonstrate the benefit of including the AGOP features to improve feature-less predictors [28, 29, 30]. The AGOP has also been show to capture surprising phenomena of neural networks beyond low-rank feature learning including deep neural collapse [31].
@@ 2.2 Alignment decomposition
$K^{(\ell)}\equiv\frac{\partial f(X)}{\partial h_{\ell}}^{\top}\frac{\partial f(X)}{\partial h_{\ell}}~{}.$ (PTK feature covariance)
@@
This alignment holds trivially and exactly if $K^{(\ell)}$ is the identity. However, the correlation can be high in trained networks where $K^{(\ell)}$ is non-trivial. For example, in the chain monomial task (Figure 1), $K^{(0)}$ is far from identity (standard deviation of its eigenvalues is $5.9$ times its average eigenvalue), but the NFA correlation is $0.93$ at the end of training. We also note that if $K^{(\ell)}$ is independent of $W^{(\ell)}$, the alignment is lower than in trained networks; in the same example, replacing $K^{(0)}$ with a matrix $Q$ with equal spectrum but random eigenvectors greatly reduces the correlation to $0.53$ and qualitatively disrupts the structure relative to the NFM (Figure 1, rightmost column). We show the same result for the CelebA dataset (see Appendix L). Therefore, the NFA is a consequence of alignment between the left eigenvectors of $W^{(\ell)}$ and $K^{(\ell)}$ in addition to spectral considerations.
Note that $K^{(\ell)}$ itself is the AGOP of the neural network $f$ with respect to the pre-activations at layer $\ell$, i.e. $K^{(\ell)}\equiv\frac{1}{n}\sum_{\alpha=1}^{n}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial h_{\ell}}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial h_{\ell}}^{\top}$. For convolutional and attention layers, the $K^{(\ell)}$ are computed by additionally averaged over all patches and token positions in the input, respectively (see Radhakrishnan et al. [21], Beaglehole et al. [22] for the formulation of the NFA in these architectures).
@@ 3 Centering the NFC isolates weight-PTK alignment
We showed that the neural feature ansatz is equivalent to PTK-weight alignment (Proposition 2). We now ask: is the increase in the NFC due to alignment of the weight matrices to the current PTK, or the alignment of the PTK to the current weights? In practice, both effects matter, but numerical evidence suggests that changes in the PTK do not drive the early dynamics of the NFC (Appendix C). Instead, we observe that the alignment between the weights and the PTK feature covariance is driven by a centered NFC that captures alignment between the parameter changes and the PTK. We then show this centered NFC can hold robustly in settings where the NFC holds with correlation less than 1, such as early in training and/or with large initialization. Finally, we establish analytically through the centered NFC how the NFA emerges as a structural property of gradient descent.
@@
where we omitted $\ell$ from all terms for ease of notation. The first term in each decomposition isolates the gram matrix of the changes in the weight matrix. We call $\bar{W}^{\top}\bar{W}$ the centered NFM and $\bar{W}^{\top}K\bar{W}$ the centered AGOP. The centered AGOP in particular measures the alignment of weight updates with the current PTK feature covariance. Both terms are 00 at initialization, and if the weight matrices change significantly (that is, if $||\bar{W}||\gg||W_{0}||$), both the NFM and AGOP are dominated by the centered terms.
@@
The centered NFC is consistently higher than the uncentered NFC across training times, architectures, and datasets Figure 2. This especially holds at early times and in deeper layers of a network (Figure 2C and D, VGG-11 on CIFAR-10 and the GPT model on Shakespeare respectively). For MLPs, we conduct a broader set of experiments and verify that the trends hold on a wide range of vision datasets. We also note that the C-NFC at early times is relatively robust to the initialization statistics of the weight matrix $W^{(\ell)}$ unlike the UC-NFC (for additional experiments, see Appendix G). We will return to this point in Section 5. These experiments suggest the C-NFC is responsible for improvements in the uncentered correlation.
High C-NFC values can drive increases in the UC-NFC as long as the centered NFM and centered AGOP are increasing in magnitude during training. We can confirm that the weights move significantly from initialization, which drives the contribution from the centered NFM and centered AGOP (Figures 14, and 15 in Appendix L). The increased importance of the C-NFC leads the UC-NFC to converge to the C-NFC at late times in most of our experimental settings, as the contribution from the weight changes dominates the contribution from the initialization. These findings validate that the C-NFC is an important contributor to increases in the UC-NFC.
Our experiments suggest that studying the C-NFC is a useful first step to understanding the neural feature ansatz. In the remainder of this section, we develop a theoretical analysis to understand why the C-NFC is generically large at early times. We then use this analysis to motivate interventions which can keep the C-NFC high and promote the NFA earlier on in training (Section 5).
3.1 Analysis at early times
@@
We note that $\bar{W}=0$ at initialization. This gives us the following proposition, which shows that early time dynamics of the centered NFM and AGOP are dominated by the second time derivatives:
@@
Let $W$ be the weights of a fully-connected layer of a neural network at initialization and $X$ be the inputs to that layer. Then, when the neural network is trained by gradient flow on a loss function $\mathcal{L}$, we have $\bar{W}^{\top}\bar{W}=\bar{W}^{\top}K\bar{W}=0$, $\frac{\operatorname{d}}{\operatorname{d}t}(\bar{W}^{\top}\bar{W})=\frac{\operatorname{d}}{\operatorname{d}t}(\bar{W}^{\top}K\bar{W})=0$, and the first non-trivial time derivatives satisfy,
@@
We immediately see from these equations how gradient-based training can drive the C-NFC to large (close to $1$) initial values - the first non-trivial time derivatives are highly correlated. This also sheds light on another mystery: namely, that the NFC does not explicitly involve the loss, yet it holds after gradient based training on the loss. The C-NFC is driven by the fact that the dynamics of the centered NFM and AGOP depend on the labels in a similar way, differing only by a single factor of the kernel $\mathcal{K}$. This eventually drives the UC-NFC and causes the NFA to hold.
When $\mathcal{K}$ is proportional to a projection matrix, i.e. $\mathcal{K}\propto\mathcal{K}^{2}$, these derivatives will be perfectly correlated. Even in the general case where $\mathcal{K}$ is dispersed with moderate eigenvalue decay, and does not commute with ˙$\dot{\mathcal{L}}XX^{\top}\dot{\mathcal{L}}$, the two quantities have non-negative covariance. In the next section we will construct a dataset which interpolates between adversarial and aligned eigenstructure to demonstrate the range of possible values the derivative correlations can take.
In the remainder of this work, we first analyze the C-NFC in early training and show that gradient based training drives the C-NFC, with a strength that depends on model and datasets (Section 4). We then discuss initializations and optimizers which can increase the C-NFC and induce larger correlation between $F$ and $\bar{G}$ (Section 5).
4 Theoretically predicting the centered NFC
In this section, we show how our decomposition for the time derivatives of the centered NFC in early times can be used to predict its value theoretically.
4.1 Early time C-NFC dynamics

### v4

@@ Abstract
Understanding the mechanisms through which neural networks extract statistics from input-label pairs through feature learning is one of the most important unsolved problems in supervised learning. Prior works demonstrated that the gram matrices of the weights (the neural feature matrices, NFM) and the average gradient outer products (AGOP) become correlated during training, in a statement known as the neural feature ansatz (NFA). Through the NFA, the authors introduce mapping with the AGOP as a general mechanism for neural feature learning. However, these works do not provide a theoretical explanation for this correlation or its origins. In this work, we further clarify the nature of this correlation, and explain its emergence. We show that this correlation is equivalent to alignment between the left singular structure of the weight matrices and the newly defined pre-activation tangent features at each layer. We further establish that the alignment is driven by the interaction of weight changes induced by SGD with the pre-activation features, and analyze the resulting dynamics analytically at early times in terms of simple statistics of the inputs and labels. We prove the derivative alignment occurs almost surely in specific high dimensional settings. Finally, we introduce a simple optimization rule motivated by our analysis of the centered correlation which dramatically increases the NFA correlations at any given layer and improves the quality of features learned.
**footnotetext: Corresponding author.$\dagger$\$$\dagger$\$footnotetext: Work partially done as an intern at Google DeepMind.
@@ 1 Introduction
Neural networks have emerged as the state-of-the-art machine learning methods for seemingly complex tasks, such as language generation (Brown et al., 2020), image classification (Krizhevsky et al., 2012), and visual rendering (Mildenhall et al., 2021). The precise reasons why neural networks generalize well have been the subject of intensive exploration, beginning with the observation that standard generalization bounds from statistical learning theory fall short of explaining their performance (Zhang et al., 2021).
@@
A promising line of work emerged in the form of the neural tangent kernel, connecting neural networks to kernels in the wide limit (Jacot et al., 2018, Chizat et al., 2019). However, subsequent research showed that the success of neural networks relies critically on aspects of learning which are absent in kernel approximations (Ghorbani et al., 2019, Allen-Zhu and Li, 2019, Yehudai and Shamir, 2019, Li et al., 2020, Refinetti et al., 2021). Other work showed that low width suffices for gradient descent to achieve arbitrarily small test error (Ji and Zhu, 2020), further refuting the idea that extremely wide networks are necessary.
@@
Subsequently, the success of neural networks has been largely attributed to feature learning - the ability of neural networks to learn representations of data which are useful for downstream tasks. However, the specific mechanism through which features are learned is an important unsolved problem in deep learning theory. A number of works have studied the abilities of neural networks to learn features in structured settings (Abbe et al., 2022, Ba et al., 2022, Nichani et al., 2023, Barak et al., 2022, Damian et al., 2022, Moniri et al., 2023, Parkinson et al., 2023). Some of that work proves strict separation in terms of sample complexity between neural networks trained with stochastic gradient descent and kernels (Mousavi-Hosseini et al., 2022).
@@
The work above studies simple structure, such as learning from low-rank data or functions that are hierarchical compositions of simple elements. Recent work makes a big step towards generalizing these assumptions by proposing the neural feature ansatz (NFA) (Radhakrishnan et al., 2024a, Beaglehole et al., 2023), a general structure that emerges in the weights of trained neural networks. The NFA states that the gram matrix of the weights at a given layer (known as the neural feature matrix or NFM) is aligned with the average gradient outer product (AGOP) of the network with respect to the input to that layer. In particular, the NFM and AGOP are highly correlated in all layers of trained neural networks of general architectures, including practical models such as VGG (Simonyan and Zisserman, 2014), vision transformers (Dosovitskiy et al., 2021), and GPT-family models (Brown et al., 2020).
@@
A major missing element is an explanation for how and why the AGOP and NFM become correlated through training with gradient descent. In this paper, we precisely describe the emergence of this correlation. We establish that the NFA is equivalent to alignment between the left singular structure of the weight matrices and the uncentered covariance of the pre-activation tangent kernel (PTK) (Section 2) features. We then introduce the centered neural feature correlation (C-NFC) which isolates this alignment process. We show empirically that the C-NFC is close to its maximum value of $1$ at early times, and fully captures the NFA at late times for a variety of architectures (fully-connected, convolutional, and attention layers) over a diverse collection of datasets (Section 3). Our experiments suggest that the C-NFC drives the development of the NFA. Through this centering, we show that the dynamics of the C-NFC can be understood analytically in terms of the statistics of the data and labels at early times (Section 4). Using this decomposition, we show that the NFA emerges as a structural property of gradient descent (analytical result in the commonly studied setting of uniform data on the sphere, under certain assumptions on the activation and target functions). In particular, the first non-zero derivatives of the centered NFM and AGOP will be asymptotically identical. We further characterize how the NFA depends on the data distribution, and explore this analytically and experimentally.
@@ 2.1 Preliminaries
One can define two objects associated with neural networks that capture learned structure. For a given layer $\ell$, the neural feature matrix (NFM) $F_{\ell}$ is the gram matrix of the columns of the weight matrix $W^{(\ell)}$, i.e. $F_{\ell}\equiv(W^{(\ell)})^{\top}W^{(\ell)}$. $F_{\ell}$ depends on the right singular vectors (and corresponding singular values) of $W^{(\ell)}$. The second fundamental object we consider is the average gradient outer product (AGOP) $\bar{G}_{\ell}$, defined as $\bar{G}_{\ell}\equiv\frac{1}{n}\sum_{\alpha=1}^{n}\frac{\partial f(x^{(\alpha)})}{\partial x_{\ell}}\frac{\partial f(x^{(\alpha)})}{\partial x_{\ell}}^{\top}$.
@@
Ansatz 1 (Neural Feature Ansatz (Radhakrishnan et al., 2024a)).
@@
The relation between the NFM and the AGOP is significant, in part, because for a neural network that has learned enough information about the target function, the AGOP of this model with respect to the first-layer inputs will approximate the expected gradient outer product (EGOP) of the target function (Yuan et al., 2023). In particular, the EGOP of the target function contains task-specific structure that is completely independent of the model used to estimate it. Where the labels are generated from a particular target function $y(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$ on data sampled from a distribution $\mu$, the EGOP is defined as
@@
We verify for this task that the AGOP of the trained model resembles the EGOP (first and third panels of Figure 1). Here the NFA holds and the NFM (second panel) resembles the AGOP and therefore the EGOP as well. Therefore, the neural network has learned the model-independent and task-specific structure of the chain-monomial task in the right singular values and vectors of the first layer weight matrix, as these are determined by the NFM. In fact, previous works have demonstrated that the NFM of the first layer of a well-trained neural network is highly correlated with the AGOP of a fixed kernel method trained on the same dataset (Radhakrishnan et al., 2024a).
This insight has inspired iterative kernel methods which can match the performance of fully-connected networks (Radhakrishnan et al., 2024a; b, Aristoff et al., 2024) and improve over fixed convolutional kernels (Beaglehole et al., 2023). Additional prior works demonstrate the benefit of including the AGOP features to improve feature-less predictors (Hristache et al., 2001, Trivedi et al., 2014, Kpotufe et al., 2016). Additionally, because the NFM is correlated with the AGOP, the AGOP can be used to recover the features from feature-less methods, such as kernel machines. The AGOP has also been show to capture surprising phenomena of neural networks beyond low-rank feature learning including deep neural collapse (Beaglehole et al., 2024).
@@ 2.2 Alignment decomposition
$\phantom{PTKfeaturecovariance}K^{(\ell)}\equiv$ $\frac{\partial f(X)}{\partial h_{\ell}}^{\top}\frac{\partial f(X)}{\partial h_{\ell}}~{}.$ (PTK feature covariance)
@@
This alignment holds trivially and exactly if $K^{(\ell)}$ is the identity. However, the correlation can be high in trained networks where $K^{(\ell)}$ is non-trivial. For example, in the chain monomial task (Figure 1), $K^{(0)}$ is far from identity (standard deviation of its eigenvalues is $5.9$ times its average eigenvalue), but the NFA correlation is $0.93$ at the end of training. We also note that if $K^{(\ell)}$ is independent of $W^{(\ell)}$, the alignment is lower than in trained networks; in the same example, replacing $K^{(0)}$ with a matrix $Q$ with equal spectrum but random eigenvectors greatly reduces the correlation to $0.53$ and qualitatively disrupts the structure relative to the NFM (Figure 1, rightmost column). We show the same result for the CelebA dataset (see Appendix M). Therefore, the NFA is a consequence of alignment between the left eigenvectors of $W^{(\ell)}$ and $K^{(\ell)}$ in addition to spectral considerations.
Note that $K^{(\ell)}$ itself is the AGOP of the neural network $f$ with respect to the pre-activations at layer $\ell$, i.e.
$K^{(\ell)}\equiv\frac{1}{n}\sum_{\alpha=1}^{n}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial h_{\ell}}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial h_{\ell}}^{\top}.$ (3)
For convolutional and attention layers, the $K^{(\ell)}$ are computed by additionally averaged over all patches and token positions in the input, respectively (see Radhakrishnan et al. (2024a), Beaglehole et al. (2023) for the formulation of the NFA in these architectures).
@@ 3 Centering the NFC isolates weight-PTK alignment
We showed that the neural feature ansatz is equivalent to PTK-weight alignment (Proposition 2). We now ask: is the increase in the NFC due to alignment of the weight matrices to the current PTK, or the alignment of the PTK to the current weights? In practice, both effects matter, but numerical evidence suggests that changes in the PTK do not drive the early dynamics of the NFC (Appendix D). Instead, we observe that the alignment between the weights and the PTK feature covariance is driven by a centered NFC that captures alignment between the parameter changes and the PTK. We then show this centered NFC can hold robustly in settings where the NFC holds with correlation less than 1, such as early in training and/or with large initialization. Finally, we establish analytically through the centered NFC how the NFA emerges as a structural property of gradient descent.
@@
where we omitted $\ell$ from all terms for ease of notation. The first term in each decomposition isolates the gram matrix of the changes in the weight matrix. We call $\bar{W}^{\top}\bar{W}$ the centered NFM and $\bar{W}^{\top}K\bar{W}$ the centered AGOP. The centered AGOP in particular measures the alignment of weight updates with the current PTK feature covariance. Both terms are 00 at initialization, and if the weight matrices change significantly (that is, if $||\bar{W}||\gg||W_{0}||$), both the NFM and AGOP are dominated by the centered terms. (Note: in the limit that initialization scale vanishes to 00 (i.e. small initialization), then we expect the centered and uncentered NFM and AGOP to coincide, as $\bar{W}=W_{t}-W_{0}\rightarrow W_{t}$, where $t$ is a step at the end of training.)
@@
The centered NFC is consistently higher than the uncentered NFC across training times, architectures, and datasets (Figure 2). This especially holds at early times and in deeper layers of a network (Figure 2C and D, VGG-11 on CIFAR-10 and the GPT model on Shakespeare respectively). For MLPs, we conduct a broader set of experiments and verify that the trends hold on a wide range of vision datasets. We also note that the C-NFC at early times is relatively robust to the initialization statistics of the weight matrix $W^{(\ell)}$ unlike the UC-NFC (for additional experiments, see Appendix H). We will return to this point in Section 5. These experiments suggest the C-NFC is responsible for improvements in the uncentered correlation.
High C-NFC values can drive increases in the UC-NFC as long as the centered NFM and centered AGOP are increasing in magnitude during training. We can confirm that the weights move significantly from initialization, which drives the contribution from the centered NFM and centered AGOP (Figures 14, and 15 in Appendix M). The increased importance of the C-NFC leads the UC-NFC to converge to the C-NFC at late times in most of our experimental settings, as the contribution from the weight changes dominates the contribution from the initialization. These findings validate that the C-NFC is an important contributor to increases in the UC-NFC.
Our experiments suggest that studying the C-NFC is a useful first step to understanding the neural feature ansatz. In Section 4, we develop a theoretical analysis to understand why the C-NFC is generically large at early times. We then use this analysis to predict the value of the C-NFC and motivate interventions which can keep the C-NFC high and promote the NFA earlier on in training (Section 5).
4 Theoretical analysis of the C-NFC at early times
4.1 Gradient flow dynamics
@@
We note that $\bar{W}=0$. This gives us the following proposition, which shows that early time dynamics of the centered NFM and AGOP are dominated by the second time derivatives:
@@
Let $W$ be the weights of a fully-connected layer of a neural network at initialization and $X$ be the inputs to that layer. Then, when the neural network is trained by gradient flow on a loss function $\mathcal{L}$, we have $\bar{W}^{\top}\bar{W}=\bar{W}^{\top}K\bar{W}=0$, $\frac{\operatorname{d}}{\operatorname{d}t}(\bar{W}^{\top}\bar{W})=\frac{\operatorname{d}}{\operatorname{d}t}(\bar{W}^{\top}K\bar{W})=0$, and the first non-zero time derivatives satisfy,
@@
We immediately see how gradient-based training can drive the C-NFC towards $1$ at early times; the first non-trivial time derivatives are often highly correlated as they differ by only a single matrix power of $\mathcal{K}$, and have the same dependence on the labels. If $\mathcal{K}$ is proportional to a projection matrix, then the two derivatives will have perfect correlation; even if not, $\mathcal{K}$ and $\mathcal{K}^{2}$ will have identical eigenvectors, and hence we might expect a range of spectra will enable high correlation between them. In the case that $\mathcal{K}$ and $\mathcal{K}^{2}$ are sufficiently correlated, the C-NFC is driven to a high value immediately upon training, and the high value of the C-NFC eventually drives the UC-NFC and causes the NFA to hold.
In the remainder of this section, we study this correlation in two different high-dimensional limits. Our analysis suggests that for large models the C-NFC has a generic tendency to increase early in training.
4.2 Maximum C-NFC for data on the sphere
We first provide a general and well-studied setting where the first non-zero derivatives of the centered NFM and AGOP are perfectly correlated - uniform data on the sphere in high dimensions (Ghorbani et al., 2020; 2021, Misiakiewicz, 2022). This limit corresponds to an infinite width network which, combined with the data symmetry, induces the PTK matrix $\mathcal{K}$ to act like a projection matrix.
Data setup
We sample data $x\sim\sqrt{d}\cdot\mathrm{Unif}\left(\mathbb{S}^{d-1}\right)$ uniformly distributed on the sphere in $d$ dimensions with radius $\sqrt{d}$. We assume the labels are generated from a target function that maps $f^{*}:\mathbb{S}^{d-1}\rightarrow[-d,d]$. I.e. the label for the point $x$ is equal to $f^{*}(x)$. We train the parameters $a,W$ for a one-hidden layer fully-connected neural network $f(x)=f(x;a,W)=a^{\top}\phi(Wx)$ with element-wise activation function $\phi$. For a learning trajectory $(a_{t},W_{t})$, the initial values $a_{0}$ and $W_{0}$ are sampled i.i.d. standard Gaussian with variance at most $O(1/k)$. I.e. for all $i,j,\ell$, we have $a_{0}(i),W_{0}(j,\ell)\sim\mathcal{N}(0,c)$ for $c=O(1/k)$.
Notation
[…]
We make the assumption that $f^{*}$ has is bounded by a polylogarithmic factor of dimension with superlinearly vanishing probability:
Assumption 4 (Labels bounded).
We have, for any positive $\delta>0$, with probability $1-O_{d}(d^{-1-\delta})$, $|f^{*}(x)|=O_{d}(\log^{\ell}(d))$ for a constant integer $\ell$.
We also assume that the target function has a non-vanishing linear component:
Assumption 5 (Non-trivial linear component).
We have, $\|\mathbb{E}_{x}\left[f^{*}(x)x\right]\|=\Omega_{d}(1)$.
These assumptions hold for example when $f^{*}$ is exactly linear, i.e. $f^{*}(x)=\beta^{\top}x$ for $\|\beta\|\leq 1$. In this case, the inner product $x^{\top}\beta=\widetilde{O}_{d}(1)$ with probability $O_{d}(d^{-1-\delta})$ for $x$ uniform on $\sqrt{d}\cdot\mathbb{S}^{d-1}$ and any positive $\delta$.
We make an additional assumption on the differentiability of the PTK, inherited from the activation $\phi$, according to the conditions from Misiakiewicz (2022). This assumption is needed for our analysis, in which differentiability enables us to Taylor expand the PTK matrix (El Karoui, 2008). We restate this assumption.
Assumption 6 (Differentiability of the PTK).
Let $\mathcal{K}(x,x^{\prime})=\mathbb{E}_{a_{k},w_{k}}\left[a_{k}^{2}\phi^{\prime}(w_{k}^{\top}x)\phi^{\prime}(w_{k}^{\top}x^{\prime})\right]=h_{d}(\langle x,x^{\prime}\rangle/d)$ be the PTK for the network $f$, where $h_{d}:[-1,1]\rightarrow\mathbb{R}$ is a positive semi-definite kernel function, defined for each dimension $d$. We assume there exist finite $h(0),h^{\prime}(0),h^{\prime\prime}(0)>0$ such that $\lim_{d\rightarrow\infty}h_{d}(0)=h(0)$, $\lim_{d\rightarrow\infty}h^{\prime}_{d}(0)=h^{\prime}(0)$, and $\lim_{d\rightarrow\infty}h^{\prime\prime}_{d}(0)=h^{\prime\prime}(0)$, where the first and second derivatives of $h_{d}$, $h^{\prime}_{d}$ and $h^{\prime\prime}_{d}$, are assumed to exist on $[-1,1]$ for all $d$.
Note this is a sub-case of Assumption 1 in Misiakiewicz (2022) for level $\ell=1$, and is satisfied for $h_{d}$ that is twice differentiable everywhere. We now introduce our final simplifying assumption on the activation function and data: